Louisiana State University Health Sciences Center Administration & Finance

Wednesday, June 27, 2007

Dental School :

  • The contractor installing the new ceiling grid and tile Work continued working on the 2nd through the 7th floors of the Clinic Building and on the 3rd floor of the Administration Building.
    • Ventilation connections are being made on the 3rd and 4th floors of the Administration Building.
  • The area clearance samples were taken on the 1st and 2nd floors of the Administration Building following the completion of the re-cleaning. 
    • Area and duct clearance samples were also taken on the 8th floor of the Clinic Building.
  • The coring that is required for the new network/telecom closets has been completed on the 2nd through the 7th floors of the Clinic Building.
    • Wire termination has begun on the 3rd floor of the Clinic building.
  • The LSUHSC carpenters completed the work to protect the antique dental equipment in the 3rd floor Administration Building breezeway.
  • LSUHSC Facility Services removed and disposed of dead trees and damaged picnic tables throughout the campus.

1542 Tulane Ave.:

  • Arc Mechanical set auxiliary air compressor #2 in place and they are making pipe connections to the unit.
    • Ordes installed the disconnect and wiring for auxiliary air compressor #2 and they are installing the wiring and receptacle box for the auxiliary air dryer #2.
  • Stratos Elevator delivered some repair parts needed for Elevator #2. 
    • They continue to make a materials list for Elevator #2  and additiona parts and materials scheduled to arrive on Thursday.

Landscaping Project:

  • All of the trees and shrubs have been planted around the CSRB. 
  • Planting in the memorial bench area has started. 
  • The installation of garden edging, the final grading and planting continues around the Lions/LSU Clinic.

Parking Lots:

  • The first of two layers of asphalt has been laid over Student Lot #3. 
    • The final layer is scheduled to be laid tomorrow.
  • The grading of Student Lot #2 is progressing.